OpenStudy (dontknowdontcare):

3. There are two fruit trees located at (3,0) and (−3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (dontknowdontcare):

I'm really stuck here, anyone help?

OpenStudy (dontknowdontcare):

@ganeshie8

OpenStudy (dontknowdontcare):

@hartnn @Mehek14

OpenStudy (dontknowdontcare):

@RhondaSommer

OpenStudy (dontknowdontcare):

@welshfella can you help?

OpenStudy (welshfella):

its been a while.... the equation of this ellipse will be x^2 / a^2 + y^2 / b^2 = 1 the center will be at (0,0) and the value of a will be 3 ( the distance from center to where graph cuts y axis)

OpenStudy (welshfella):

so we have x^2 y^2 --- + --- = 1 9 b^2

OpenStudy (welshfella):

sorry gtg right now I hope someone else can help you

OpenStudy (dontknowdontcare):

@retirEEd can you help?

OpenStudy (retireed):

Sorry I was off researching how to do this and I am unfortunately still not sure how to do this question.

OpenStudy (dontknowdontcare):

Yeah, Im really stuck on it Thanks for trying though.

OpenStudy (dontknowdontcare):

@mrm sorry to bother but do you know this?

OpenStudy (mrm):

The a and b in your ellipse equation come from your points. So (-3,0) and (3,0) are equivalent to \[{(-\sqrt{a^2-b^2}, 0)\space and\space (\sqrt{a^2-b^2}, 0)}\]

OpenStudy (mrm):

This however is not enough information to create an ellipse. The instructions require you to create two vertices on the y-axis. The vertices should each be equal distance from the origin (0,0) so that the center of the ellipse is at (0,0). So choose two points on the y-axis to create your vertices. Do that now.

OpenStudy (mrm):

You can choose any two points you want.

OpenStudy (mrm):

Once you have your vertices (two points on the y axis) and you have you two foci (given) You are equipped with all the information you need to find a and b fr you ellipse equation: \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] a is your major axis (the longest axis of the ellipse b is your minor axis (the shortest axis of the ellipse)  The equation for your hyperbola is \[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]